OpenStudy (anonymous):

an empty bucket is put under two faucets. if one faucet is turned on alone, the bucket fills in 6 minutes.if the other faucet is turned on, it fills in 4 minutes. if both are turned on, how many seconds will it take to fill the bucket?

OpenStudy (anonymous):

Let x be the answer Make this chart: buckets filled time rate in buckets/minute 1st faucet 1 6 2nd faucet 1 4 both open 1 x Fill in the rates by dividing buckets by minutes: buckets filled time rate in buckets/minute 1st faucet 1 6 1/6 2nd faucet 1 4 1/4 both faucets 1 x 1/x Equation: 1st faucet's rate + 2nd faucet's rate = their combined rate: 1/6 + 1/4 = 1/x Multiply through by 12x 2x + 3x = 12 5x = 12 x = 2.4 minutes.